EXPERIMENTAL DETERMINATION OF THE ENERGETIC LOCATION OF THE LOWEST LYING ELEC-TRONICALLY EXCITED STATE OF H_{3}O RYDBERG RADICAL

\underline{MICHAEL H. COHEN}, WAN YONG FENG, ELLEN GUNDLACH, and JAMES V. COE, Department of Chemistry, The Ohio State University, Columbus, OH 43210-1173..

The lowest lying electronic state of H_{3}O Rydberg radical is determined to lie 2.3 eV above the ground state. This result has been obtained by studying the reaction H_{3}O^{+} + OH^{−} in the gas phase as a function of center-of-mass collision energy using electrostatically merged ion beams. The rate constant of this reaction was measured as a function of center-of-mass collision energy and put onto an absolute scale by calibrating it to the reaction O^{+} + O^{−} and using Landau-Zener (LZ) absorbing sphere theory. When applying LZ absorbing sphere theory to ion-ion recombination reactions, one important parameter is the location of electronically excited states of the neutral products. It is this dependence that has allowed an experimental determination of the location of the lowest lying electronic state of H_{3}O Rydberg radical.